The sum of two numbers is 60, and the sum of their reciporcals is 5/72. What are the numbers?
Answer by stanbon(75887)

The sum of two numbers is 60, and the sum of their reciprocals is 5/72. What are the numbers?
----------------------
Equations:
x + y = 60
1/x + 1/y = 5/72
-----
x = 60-y
----
Substitute for "x" and solve for "y":
1/(60-y) + 1/y = 5/72
---
72y + 72(60-y) = 5y(60-y)
---
72*60 = 300y-5y^2
----------------
5y^2-300y+72*60 = 0
---
y^2-60y+12*72 = 0
(y-36)(y-24) = 0
y = 36 or y = 24
========================
If y = 36, x = 24
If y = 24, x = 36
--------------------------
cheers,
stan H.